# Learnable Fractal Flames

### v1

@@ 3.1. Learning Non-Linear Flames
In Scott Draves’ description of the fractal flame algorithm (Draves and Reckase, 2008), 48 different non-linear generator function variations were introduced. Each of these variations can be combined with others in an infinite number of different ways. In our formulation, each generator need only be a differentiable function $f:X\rightarrow X$, which leaves substantial room for experimentation and exploration. Accepting that a comprehensive study of this space is impossible, we first demonstrate our approach using a small selection of non-linear variations.
Figure 4 shows a reference image alongside optimized fractals representing a linear variation and 6 non-linear variations. Many of these fractals are not capable of capturing the structure of the reference image closely. This is especially true when the intrinsic structure of the fractal does not match the reference image. Nevertheless, they all show a tendency to assign color and density to appropriate areas of the image. In these experiments, the background color of the image was set to white and was not learned. Each fractal was trained at a resolution of 100x100 pixels and has 8 generator functions.
3.2. Learning Multi-Flame Art
Next, we trained multi-flame fractals to emulate two famous paintings in the public domain (”Fighting Forms” by Franz Marc (Marc, 1914), and ”Wisteria” by Claude Monet (Monet, 1925)). The results are shown in Figure 1. For Wisteria, we composed four layers of linear fractal flames with 24 generator functions each. For Fighting Forms, we composed 3 layers of non-linear fractal flames including spherical, hankerchief, and exponential variations. Both fractals were trained at a resolution of 200x200 pixels. In this experiment, we found that our pipeline was capable of capturing the gross structure of the reference image quite closely, while still exhibiting fractal aesthetic qualities.
3.3. Exploring Hyper-Parameter Effects
Figure 5. Two fractal flames that were trained on the same reference images as Figure 1 but with modified hyper-parameters. The left fractal was trained at a resolution of 50x50 instead of 200x200. The right fractal was trained with 8 generator functions per flame instead of 24 generator functions.
@@ 3.3. Exploring Hyper-Parameter Effects
Two fractal flames that were trained on the same reference images as Figure 1 but with modified hyper-parameters.
Figure 6. Two linear fractal flames that were trained on the same reference image as Figure 4 but with different hyper-parameters. The left fractal was trained at a resolution of 20x20 instead of 100x100. The right fractal was trained with 3 generator functions instead of 8.
[…]
Two linear fractal flames that were trained on the same reference image as Figure 4 but with different hyper-parameters.
Next, we explored the effects associated with altering the training image resolution. The non-linear Fighting Forms fractal experiment was re-created using a training image resolution of 50x50 pixels instead of 200x200 pixels (Figure 5). The effect was to produce a fractal with significantly more thin and sharp structures compared to the original experiment. Because samples are grouped together in larger pixels at lower resolutions, the optimization process does not encourage samples to spread out to the same degree. Figure 6 replicates this effect in a linear fractal trained on the same simple reference image as in Figure 4 using a resolution of 20x20 pixels instead of 100x100 pixels.
We also explored the effects associated with altering the number of generator functions in a fractal flame. The linear Wisteria fractal experiment was re-created using 8 generator functions for each flame instead of 24 (Figure 5). The effect was to produce a fractal with simpler structure and color palette. This effect is quite intuitive as reducing the number of generator functions reduces the number of parameters in the sampler, as well as the number of color parameters in the painter. Figure 6 replicates this effect in a linear fractal trained on the same simple reference image as in Figure 4 using 3 generator functions instead of 8. Interestingly, the simpler fractal converged to a warped version of the Sierpinski triangle.
Finally, we explored the use of different generator function variations in a multi-flame pipeline. Figure 7 shows fractal images trained on Fighting Forms and Wisteria using different non-linear variations. Each image reflects the gross structure of the reference image while retaining the fractal aesthetics of the different variations. These experiments used a composition of 3 fractal flames with 24 functions each, trained at a resolution of 200x200 pixels
Figure 7. Fractal flames that were trained on the same reference images as Figure 1 but with different generator function variations. Top left: linear variation. Top right: spherical variation. Bottom left: fisheye variation. Bottom right: power variation.
@@ 4. Limitations
The differentiable fractal rendering approach described here has several limitations. First, the loss landscape associated with learning IFS fractal parameters is highly non-convex. This means that that the initial configuration of parameters has a strong effect on the final result. While we were able to learn interesting and aesthetic fractals from random initialization, we occasionally needed to play with the randomization settings or random seed before we arrived at a satisfactory result. For this reason, we believe that combining a differentiable rendering approach with the manual editing of fractal parameters will likely be the approach with the most artistic utility.
Another limitation of our approach is the potential for training instability when using non-linear variations. In practice, we found that this was rarely an issue. However, we did not explore all possible variations and stability issues may still be encountered in some scenarios.
@@ 5. Conclusion and Future Work
In this work, we extended the state-of-the-art differentiable IFS fractal rendering approach to support non-linear generator functions, color images, and multi-flame compositions. Our experiments show that differentiable rendering is now a viable approach for creating fractal art. Overall, we believe that combining differentiable rendering with generative simulations is a fascinating area of research that can empower artists with new creative tools. Future work may explore higher dimensional IFS fractals, other fractal varieties, and other types of generative simulations. We hope that our work will inspire others to explore this space as well.

### v2

@@ 3.2. Exploring the Learning Process
Another important aspect of the learning algorithm is a high degree of sensitivity to initial conditions. This means that choosing different initial values for the fractal parameters (prior to running the optimization process) results in different final fractal parameters and visually different results. Figure 6 shows several outputs from fractal optimization processes that differed only in that they used different random seeds to initialize the fractal parameters. These results indicate that the loss landscape associated with the optimization process is non-convex and, therefore, that the optimization processes are finding different local minima rather that a common global minimum. This fact is highly relevant to the artistic application of this optimization process, as it means that artists may achieve different outcomes by manually configuring the initial parameters or by initializing with different random seeds.
Figure 6. Four fractal flames with 8 generator functions that were trained using different initial parameter values. When the initial parameters are different, the optimization process will often find a different local minimum, rather that a common global minimum. All fractals are linear and were trained on the reference image shown in Figure 4.
@@
Another interesting aspect of fractal optimization that we observed was that the resolution of the image used during training had a strong effect on the outcome of the optimization process. The splatting operation that enables the fractal rendering process to become differentiable, applies a Gaussian blur kernel with a standard deviation of 0.5 pixels to the resulting image. Because the magnitude of the blur is a function of the pixel size, reducing image resolution results in a larger blurring kernel. A larger blurring kernel affects the gradients that are used in the optimization process, which results in different outcomes. Figure 7 shows several outputs from fractal optimization processes that differed only in that they used different image resolutions when optimizing the fractals. The effect of using smaller image resolutions was generally to produce fractals with more thin and sharp structures.
Figure 7. Top: a linear fractal (left) and non-linear spherical fractal (right) trained at a resolution of $20\times 20$. Bottom: a linear fractal (left) and non-linear spherical fractal (right) trained at a resolution of $400\times 400$. Fractals trained at lower resolutions tend to have more thin and sharp structures. All fractals used 8 generator functions and were trained on the reference image shown in Figure 4.
@@
3.3. Multi-Fractal Compositions
Finally, we trained fractals with the objective of creating complex visual art. We selected two famous paintings in the public domain to use as reference images for these experiments (Fighting Forms by Franz Marc (Marc, 1914), and Wisteria by Claude Monet (Monet, 1925)). All experiments discussed in this section used a training image resolution of $200\times 200$ pixels.
Our initial experiments used a single fractal flame with 24 generator functions for each image. Some of these results are shown in Figure 8. With the goal of increasing the complexity of the generated fractal artwork, we began experimenting with multi-fractal compositions. Figure 1 shows two results from these experiments. For Wisteria, we composed 4 linear fractal flames with 24 generator functions each. For Fighting Forms, we composed 3 non-linear fractal flames, which included spherical, handkerchief, and exponential variations. In this experiment, we found that our pipeline was capable of capturing the gross structure of the reference image quite closely, while still exhibiting the desirable fractal aesthetic qualities.
Figure 8. A non-linear spherical (left) and a linear (right) fractal flame trained on the same reference images as Figure 1. These experiments used only a single fractal, unlike those in Figure 1 which composited multiple fractals.
[…]
Individual fractal flames that were trained on the same reference images as Figure 1.
@@ 4. Limitations
The differentiable fractal rendering approach described here has several limitations. One limitation of our approach is the potential for training instability when using non-linear variations. In practice, we found that this was often not an issue. However, we did not explore all possible variations and stability issues may still be encountered in some scenarios.
@@ 5. Conclusion and Future Work
In summary, we developed a state-of-the-art differentiable IFS fractal rendering approach that supports non-linear generator functions, color images, and multi-fractal compositions. Differentiable fractal rendering in combination with gradient descent optimization allow artists to use reference images in order to control the generation of colorful and complex fractal flame artwork. An exciting avenue for future work would be to integrate this approach into interactive graphical applications similar to existing fractal flame creation software (Glare Technologies, 2015; Sdobnov et al., 2009; Draves, 1999; Maschke, 2001). This would allow fractal artists without a background in programming to access and explore this technology.